OpenStudy (sepeario):

What is the smallest number that 45 can be multiplied by to give a perfect square?

OpenStudy (jennychan12):

use trial and error.

OpenStudy (sepeario):

Is there a better way?

hartnn (hartnn):

yes, there's a way.

hartnn (hartnn):

can you factor 45 ?

hartnn (hartnn):

@Sepeario ?

hartnn (hartnn):

ok, to get a perfect square, we need pairs of numbers, like 49=7*7, pair of 7. 100 = 10*10 = 2*2* 5*5 , pairs of 2 and 5. similarly, if you want a perfect square, you just have to complete squares, say for example you have 28. first you factor 28 = 2*2*7 now you see, that 2 is forming a pair, but 7 doesn't. so, the smallest number which you have to multiply to 28 is = 7 to make it perfect square, (because 2*2*7*7 is a perfect square of 2*7=14) can you do the same thing with 45 ?

OpenStudy (sepeario):

@hartnn Thanks.

hartnn (hartnn):

so which number will you multiply ?

OpenStudy (anonymous):

5

OpenStudy (shubhamsrg):

the smallest PERFECT square is 1 right ? 1^2 = 1 in the question, we are asked the smallest real no. which when multiplied by 45 yields a perfect sq, so , shouldnt ans be 1/45 ? ;) i know that shouldnt be correct, but the asked hasnt put up a decent language to the question, and so,,why not play! :D

OpenStudy (shubhamsrg):

asker*

hartnn (hartnn):

@shubhamsrg , 0 is a perfect square , isn't it ? so, smallest number---->0 smallest natural number---->5

OpenStudy (hba):

0 is a very imp number :)

OpenStudy (shubhamsrg):

ahh yep..0 !!

OpenStudy (jiteshmeghwal9):

First factor 45It needs just one 5 more to make a pair of two equal numbers so the smallest number is 5.